The third side of an isosceles triangle is 2 inches shorter than each of the other two sides. If the perimeter of this triangle is 67 inches, find the length of each side.
Answer by rfer(16322)

2x+x-2=67
3x-2=67
3x=69
x=23in
x-2=21in
21+46=67